OpenStudy (anonymous):

Math help plz asap win medal Julie solved eachof the equations. To solve the first equation, she added first and then multiplied.To solve the second equation, she multiplied first then added. Why did Julie add and multiply in a different order each time? Explain your thinking. 1/8a-3/4=1/2 a-5/9=1/3

OpenStudy (anonymous):

I's on my math book but it's called core focus i hate those can you help me?

OpenStudy (anonymous):

Well try it yourself, you have the equation 1/8a-3/4=1/2 do the steps Julie did and see what you get and see if it's right or not, it's just simple algebra really :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so for the first equation it would be 1/8a+3/4=12 that's step on because she added

OpenStudy (anonymous):

No, you want to isolate a, meaning get a by its self. \[1/8a-3/4=12 \implies 1/8a-3/4\color{red}{+3/4}=12\color{red}{+3/4}\]

OpenStudy (anonymous):

As your question doesn't specify what she is adding/ multiplying by so we will have to go with this :)

OpenStudy (anonymous):

ok 3/4+3/4 is either take 3/4 away or add it up

OpenStudy (anonymous):

but 12+3/4 would be 12 and 3/4?

OpenStudy (anonymous):

You want to get `a` by itself, so to do that we use the opposite operation to move everything other than a to the right.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12+3/4 = 51/4

OpenStudy (anonymous):

\[1/8a=51/4\] so now we have this

OpenStudy (anonymous):

ok so then 1/8a=51/4 right what we could do is 1/8 divided by 1/8 on both sides

OpenStudy (anonymous):

Nope because we want to cancel out the 1/8 on the left, if we divide by 1/8 on the left side we would just get 1/64a

OpenStudy (anonymous):

So why don't we multiply by 8 on both sides instead? :P

OpenStudy (anonymous):

\[8 \times \frac{ 1 }{ 8 }a = \frac{ 51 }{ 4 } \times 8\]

OpenStudy (anonymous):

so then next we multiply 51/4*8=102

OpenStudy (anonymous):

Yes, so a = 102

OpenStudy (anonymous):

ok thank you